# DeLTa: Demonstration and Language-Guided Novel Transparent Object Manipulation

### v1

@@ V Conclusion
We propose DeLTa, a framework that integrates novel object 6D pose estimation, precise long-horizon manipulation of transparent objects from language instructions. Our last-inch motion planner generalizes 6D trajectories to novel objects from single-object demonstrations, while our VLM-guided planner grounds task plans in robot configurations and refines them for long-horizon manipulation tasks. Our method demonstrates robust transparent manipulation capabilities in various real-world environments and challenging manipulation tasks, substantially outperforming existing competitive methods.
Limitations: DeLTa is currently limited to rigid object manipulation, as our pose estimation method assumes a rigid-body model. Also, we implemented only three primitive skills and ten primitive actions, as a proof of concept, while leaving broader task diversity and complexity for future work. Extending DeLTa to handle transparent deformable objects, such as plastic bags, and to incorporate a broader range of primitive actions across diverse objects would be promising directions for future work.

### v2

@@ V Conclusion
We propose DeLTa, a Demonstration and Language-Guided Novel Transparent Object Manipulation framework, that integrates novel object 6D pose estimation and precise long-horizon manipulation of transparent objects from language instructions. Our last-inch motion planner generalizes 6D trajectories to novel objects from single-object demonstrations, while our VLM-guided planner grounds task plans in robot configurations and refines them for long-horizon manipulation tasks. Our method demonstrates robust transparent-object manipulation capabilities in real-world environments and challenging tasks, substantially outperforming existing competitive methods.
Limitations: First, DeLTa is currently limited to rigid object manipulation, as our pose estimation method assumes a rigid-body model. Second, our current implementation relies on pre-built canonicalized meshes for cross-object trajectory transfer and supports three primitive skills and ten primitive actions as a proof of concept. Extending both the mesh preparation pipeline and task diversity would be promising directions for future work.